Let f(x, y, z)=zxly. Find the value of the following partial derivatives. (a) f(4,4,3) (b) fy(4, 2, 4) (c) f(2,4,3)
Answer #1

fx(x,y,z) = (1/y)*ln(z) * zx/y

fx(4,4,3) = (1/4) * ln(3) * 34/4 = (3/4)ln(3) = 0.823

fy(x,y,z) = ln(z) * (z)x/y * (-x/y2)  

fy(4,2,4) = ln(4)*(4)4/2 * (-4/22) = -32 ln(2) = - 22.18

  fz(x,y,z) = (x/y) * (z)(x/y)-1

  fz(2,4,3) = (2/4) * (3)(2/4)-1 = 1/2√3 = 0.288